OpenStudy (anonymous):

You are designing a marble basin that will hold a fountain at the center of a zoo. The basin's slides and bottom should be 1 foot thick. Its outer length should be twice its outer width and outer height. What should the outer dimensions of the basin be if it is to hold 36 cubic feet of water?

OpenStudy (anonymous):

are there any options???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok i know this has nothing to do with your question But your name is really familiar. ok do you know anything about JIS?

OpenStudy (anonymous):

No, is that a name or something?

OpenStudy (welshfella):

let the length of the inner part be x then outer length = x + 2 outer width will be 0.5(x + 2) outer height = 0.5(x + 2) inner width = 0.5(x + 2) - 2 = 0.5x - 1 inner height = 0.5(x + 2) - 1 = 0.5x

OpenStudy (anonymous):

Nope. theres was a girl in my class in JIS and she had the same name Lol

OpenStudy (welshfella):

so inner volume = x(0.5x)(0.5x - 1) = 36 solve for x

OpenStudy (anonymous):

@Welshfella is right. :)

OpenStudy (welshfella):

only problem is you have a cubic equation . Maybe there's an easier way...

OpenStudy (welshfella):

that one simplifies to 0.25x^3 - 0.5x^2 - 36 = 0

OpenStudy (welshfella):

you can either solve this by trial and improvement or graphically or using a graphical calculator ( thats the easiest way - lol!!)

OpenStudy (anonymous):

would the answer be 4?

OpenStudy (welshfella):

i'll check my calc i get x to be 5.369 m

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (welshfella):

oops sorry I made an error !!

OpenStudy (welshfella):

;x comes to exactly 6

OpenStudy (welshfella):

that is the internal length but the question asks for the outer dimensions so outer length = 6 + 2 = 8 outer width = 0.5 * 8 = 4 and outer height = = 4